OpenStudy (anonymous):

Rewrite in simplest rational exponent form √x • 4√x

OpenStudy (alexandervonhumboldt2):

this is easy. what is \[\sqrt{x} * \sqrt{x}\]?

OpenStudy (alexandervonhumboldt2):

there is a rule: \[\sqrt{a}*\sqrt{b}=\sqrt{a*b}\]

OpenStudy (alexandervonhumboldt2):

so square root of x multiplied by square root of x will be equal ...... square root of x squared \[\sqrt{x^2}\]